Compute P(X) using the binomial probability formula. Then determine whether the normal distribution can be used to estimate this probability. If so, approximate P(X) using the normal distribution and compare the result with the exact probability.
n=58, p=0.3, and X=24

Answer by stanbon(75887)

Compute P(X) using the binomial probability formula. Then determine whether the normal distribution can be used to estimate this probability. If so, approximate P(X) using the normal distribution and compare the result with the exact probability.
n=58, p=0.3, and X=24
p*n = 0.3*58 > 5
q*n = 0.7*58 > 5
------
P(x = 24) = P(23.5 < x < 24.5)
np = 0.3*58 = 17.4
sqrt(npq) = sqrt(17.4*0.7) = 3.49
--
z(23.5) = (23.5-17.4)/3.49 = 1.748
z(24.5) = (24.5-17.4)/3.49 = 2.034
---
P(x = 24) = P(1.748< z < 2.034) = 0.0193 (normal approximation)
--------------------
P(x = 24) = 58C24(0.3^24)(0.7^34) = 0.0196
----------------
Cheers,
Stan H.
-----------------